OpenStudy (anonymous):

So,I need help with this question. The position of a particle moving under uniform acceleration is a function of time and the acceleration. Suppose we write this position as x=kamtn, where k is a dimensionless constant. Find m and n by dimensional analysis. pls explain it as detail as possible.thx a lotttt!

OpenStudy (michele_laino):

is your function like this: \[x\left( t \right) = k \cdot a \cdot m \cdot t \cdot n\]

OpenStudy (anonymous):

Think about what the units on "a" and "t" are first, start there and see how you can get to units of just "meters" (since X is in units of just meters)